OpenStudy (anonymous):

Integrate: 2x cosh(x) dx, upper limit=ln(2), lower limit=0

OpenStudy (kropot72):

Is this it? \[\int\limits_{0}^{\ln 2}2x.\cosh x.dx\]

OpenStudy (anonymous):

You have in the integrand 2x*cosh(x) But you know that cosh(x)=(e^x + e^-x)/2=(e^x)/2 + (e^-x)/2 So you have in the integrand (what's inside the integral): 2x[(e^x)/2 + (e^-x)/2]=xe^x + xe^-x First term Second Term To integrate the first term xe^x you use integration by parts Let u=x ===> du=dx Let dv=(e^x) dx ====> v=integral of e^x so v=e^x Plug this into your formula integral of vdu= uv- integral vdu so integral of xe^x= xe^x - integral e^x dx so integral of xe^x=xe^x -e^x PART A SOLUTION Now to integrate the second term you use integration by parts again, it will be very similar. Second term is xe^-x Integral of xe^-x dx= uv - integral vdu Let u=x ====> du=dx Let dv=e^-x dx =====> v=integral of e^-x dx, so v=-e^-x Thus integral of xe^-x dx=x (-e^-x) - integral -e^-x dx =-xe^-x-e^-x PART B SOLUTION So your answer is PART A SOLUTION + PART B SOLUTION evaluated from 0 to ln2 ANSWER: = [ln2e^(ln2) -e^(ln2) -ln2e^(-ln2)-e^(-ln2)] - [0e^0 -e^0 -0e^-0 -e^-0] Now simplify.

OpenStudy (anonymous):

Thanks mashe, that helped a lot. kropot: yes, that's the question.

OpenStudy (anonymous):

I haven't integrated hyperbolic functions in a long time (but I simply used what it is in exponential form) there might be an easier way (IDK maybe there isn't)...please check if I made any mistakes...but this is pretty much the strategy. I suggest as I suggest to anyone i provide help to try it yourself after understanding the steps you have to take and see if you arrive at the correct answer.

OpenStudy (anonymous):

Oh and Of course you are welcome! More than happy to help :-)

OpenStudy (anonymous):

Hey mashe, I think there may be a slightly shorter way, cosh(x) simply integrates to sinh(x)... is this correct?

OpenStudy (anonymous):

HAHHAHA YES!!! AWESOME :-)

OpenStudy (anonymous):

I always take the longer route I like the extra work, seriously all my teachers tell me this...

OpenStudy (anonymous):

good :) but in this question they give you some extra info (hints): cosh(ln(2))=5/4, sinh(ln(2))=3/4

OpenStudy (anonymous):

Okay awesome here it goes: So then integral of 2x coshx dx= 2* Integral x coshx Let u=x =====> du=dx Let dv=coshx dx======> v=sinhx So we have 2* Integral x coshx=2[ uv - integral vdu] =2 [xsinhx- integral sinhx dx] the terms in the parenthesis are evaluated from x=0 to x=ln2 =2[xsinhx-coshx] =2 {[ ln2 sinh(ln2) - cosh(ln2)]- [0sinh(0) -cosh(0)] } = 2{[ln2 (3/4) -(5/4)]-[0 sinh(0)-cosh(0)]}

OpenStudy (anonymous):

sinh(0)=0 and cosh(0)=1 simplify :-)

OpenStudy (anonymous):

Which then equals \[(3(\ln (2))/2 - 1/2\] Is this correct?

OpenStudy (anonymous):

It would have to be correct haha I just looked up the answer and that's the right one! Thankyou mashe!

OpenStudy (anonymous):

You are welcome Saltin8R happy to help and better when it is the easier way or the most appropriate way hahaha